OpenStudy (anonymous):

really could use some help The half-life of a certain radioactive material is 32 days. An initial amount of the material has a mass of 361 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 5 days. Round your answer to the nearest thousandth.

OpenStudy (anonymous):

;0.199KG

OpenStudy (anonymous):

;323.945KG

OpenStudy (anonymous):

Lets look at what half life means: A half-life of 32 days means that after 32 days a mass of 1 will reduce to 1/2. \[e^{k\cdot 32}=\frac{ 1 }{ 2 }\] take the natural log of both sides.\[k \cdot 32=\ln (1/2)=-\ln 2\] divide by 32 on both sides.\[k=\frac{- \ln 2 }{ 32 }\]

OpenStudy (anonymous):

So this gives us our formula for the fraction of material after time "t" as:\[e^{-\frac{ \ln 2 }{ 32 } t}\]

OpenStudy (anonymous):

;0.797kg

OpenStudy (anonymous):

;0kg

OpenStudy (anonymous):

that's abcd I really don't get it at all

OpenStudy (anonymous):

abcd?

OpenStudy (anonymous):

THATS WHAT I HAVE TO CHOOSE FROM

OpenStudy (anonymous):

I don't understand it at all

OpenStudy (anonymous):

So for exponential growth or decay the equations are like this. \[(\text{initial stuff})\cdot e^{k \cdot t}=\text{(stuff after time t)}\] We just have to figure out what "k" is. "k" will be negative when it is decaying. "k" will be positive when it is growing.

OpenStudy (anonymous):

ok so what is the e for? an what about the g

OpenStudy (anonymous):

e is just the natural exponent it is a constant for all of these equations e is approximately 2.71828

OpenStudy (anonymous):

For half life situations (exponential decay) K is always as follows:\[k=-\frac{ \ln 2 }{ \text{half life} }\]

OpenStudy (anonymous):

half-life of the substance : A = 391*2^(-5/32) using a calc find 2^(-5/32) A = 391 * .89735 A = 350.866 I keep getting that an I don't no why because its not on there

OpenStudy (anonymous):

In our situation what is k equal to?

OpenStudy (anonymous):

so how do I find the answer

OpenStudy (anonymous):

I'm trying to walk you through it.

OpenStudy (anonymous):

ok so after the For half life situations (exponential decay) K is always as follows what is next

OpenStudy (anonymous):

So what class is this for? I guess maybe you aren't using real half lifes but approximate ones. So your expressions should be like this:\[\text{initial stuff}\cdot (1/2)^{\frac{ 1 }{ \text{halflife} }\cdot t}=\text{stuff after time t}\]

OpenStudy (anonymous):

yea I don't get how they wrote it but thanks for helping anyways an after that

OpenStudy (anonymous):

Sorry for the confusion.